OpenStudy (anonymous):

Sally has 12 coins consisting of quarters and dimes. The value adds to $2.25, how many of each coin does she have?

OpenStudy (anonymous):

let x be quarters and y be dimes. so x + y = 0.35 agreed?

OpenStudy (anonymous):

Where did 35 cent come from ?

OpenStudy (anonymous):

a quarter is 0.25 $ and a dime is 0.10$

OpenStudy (anonymous):

ok , ok .

OpenStudy (anonymous):

agreed.

OpenStudy (anonymous):

lol, sorry, ignore that 0.25x + 0.10(12-x) = 2.25¢

OpenStudy (anonymous):

so , when I multiply .10 by 12 would I get 1.2x ?

OpenStudy (anonymous):

yes or yu can have 25x+10(12-x)=225

OpenStudy (anonymous):

will x = 1.6 ?

OpenStudy (anonymous):

0.25x + 0.10(12-x) = 2.25¢ 0.25x+1.2-0.1x = 2.25 o.15x = 1.05 x=7 so there's 7 quarters

OpenStudy (anonymous):

and 5 dimes ?

OpenStudy (anonymous):

yea so let's check it remember i said at first one quarter plus one dime =0.35 so what is (7*0.25) + (5*0.10)

OpenStudy (anonymous):

7*0.25= 1.75 and 0.10*5=.50 1.75+.50 = 2.25